# Pay Attention: a Call to Regulate the Attention Market and Prevent Algorithmic Emotional Governance

### v1

@@ 1 An unsustainable attention market
Since the advent of mass market in the 50’s, media and advertisement providers have relentlessly tried to figure out effective methods to capture our attention and turn it into revenue. During the last two decades, supported by advances in artificial intelligence (AI), major online social media and Web platforms have brought this process of capturing attention to an unprecedented scale. Based almost exclusively on advertising revenues, their business model consists in providing free services that, in return, collect behavioral traces. This data is then used to maximize the impact of advertisements on users by 1 ensuring their mental availability at the time of being shown the advertisement, and 2 ensuring that the message meets their interests, beliefs and moods (i.e. targeted advertising). Based on research in psychology, sociology and neuroscience, several actors including online social media, games and Web platforms have engineered techniques capable of very effectively plundering our “available brain time” [37, 27]. We can distinguish two broad categories of such techniques. Firstly, some techniques are explicitly designed to leverage cognitive biases as a means to capture attention. For instance, the likes collected after posting content activate the brain’s dopaminergic pathways (involved in the reward system) and tap into our need for social approval, giving “bright dings of pseudo-pleasure” [37]; notifications of smartphone applications feed our appetite for novelty and surprise such that it is difficult to resist the urge to check them; the pull-to-refresh mechanism [37], alike slot machines, exploits the variable reward pattern whereby each time we pull down the screen we may get an update or nothing at all; infinite scrolling (of news, posts or videos…) traps us because of our fear of missing out important information (FOMO) to the point that we can hardly break the flow; automatic video chaining replaces a deliberate action to continue watching with a required action to stop watching, and entails a frustrating feeling of incompleteness when stopped; etc. Similarly, some techniques harness dark patterns [^1] The legal definition in California is “A user interface is a dark pattern if the interface has the effect of substantially subverting or impairing user autonomy, decision-making, or choice. A business’s intent in designing the interface is not determinative in whether the user interface is a dark pattern, but a factor to be considered.” CPRA § 7004 (c) [26] to manipulate users into taking actions or decisions they wouldn’t take otherwise. This is typically the case when one accepts all notifications of an application without really noticing it, while deactivating notifications would require an additional, less intuitive, series of actions.
@@
Secondly, recent advances in machine learning allow the training of content recommendation algorithms on massive online behavioral data. These algorithms learn to recommend content that not only captures attention but also increases user engagement [^2] There are multiple definitions of user engagement. In the context of social media, this typically refers to the fact that a user would interact with a content: e.g. like, comment or repost it. Engagement is usually public in that it leaves public traces on the platform, unlike sheer content consumption that remains private [50].. They discover the content’s key features that help predict whether such content will effectively attract users’ attention, and typically end up selecting content related to conflictuality, fear or sexuality [10]. They also learn to exploit humans’ negativity bias [61, 59] and, as a consequence, content conveying high-arousal negative emotions (such as anger, resentment, indignation and disgust) are more likely to be read and eventually shared online than those conveying other emotions [50, 35]. Concerningly enough, false information (a broad term including misinformation and other forms of disinformation) typically relies on such negative emotions as a trick to foster sharing. Finally, recommendation systems may do all this without it being explicit in the features they select, nor in the succinct feedback that some of them happen to provide [^3] For instance, a recommendation system may tell us “you liked this movie, you may also like this one”. But we don’t know what features were selected to recommend this one: Do they have an actor in common? Did my contacts like both of them? etc..
@@
Since the amount of attention available is both limited and precious, it would be unsustainable for a civilization to waste it with impunity for questionable or futile purposes [10]. Today, we might precisely be at that moment: while mental time has become a new oil, we have created an attention economy and subsequent attention markets [28, 29] that, although sustainable from an economic point of view, may be unsustainable from a civilization point of view. From these first references, let us define what the term “attention market” refers to in this article.
@@
The attention market treats attention as a tradable commodity and involves multiple actors: from producers (the end users whose attention is the resource), to content creators whose work is used to capture the attention, brokers who trade and monetize the attention, and consumers who use it for their purposes such as exposing users to advertisements.
In this article, we propose a discussion aimed at spurring introspection and debate within the computer science community. In line with the Web Science Manifesto [6] calling for interdisciplinary approaches to prepare the future of the Web, we bring together and synthesize the conclusions of more than 70 papers and books from a wide range of disciplines to analyze the practices and drifts of these systems designed to capture attention on a worldwide scale. We make the point that, with the initial commonplace goal of making targeted advertising more effective, the generalization of attention-capturing techniques and their use of negative emotions tends to foster radicalization and polarization, amplify the dissemination of false information, spur the emergence of populism, and eventually put a threat on democracies and human societies in general.
Promoting awareness about these issues, this paper is directly related to UN’s Sustainable Development Goal [^4] https://www.un.org/sustainabledevelopment/ (SDG) 16 “Peace, Justice, and Strong Institutions”, since it suggests actions to combat the instrumentalization of negative emotions, the associated false information that mechanically increase the level of anger and resentment among populations, and it promotes “societies that respect the right (…) to freedom of expression, and access to information” [^5] https://www.un.org/sustainabledevelopment/peace-justice/. By pointing to the rise of populism worldwide, it addresses the connected question of how to strengthen institutions. The paper is also relevant with respect to SDG 3 ”Good Health and Well-being” considering the aggravating effects of online social media on mental health, and the public health issues caused by false information (e.g. during the Covid-19 pandemics).
So far, the public and private research in computer science has invested large efforts in dealing with some aspects of the problem like radicalization, violent speech and false information. These works rely on post hoc measures such as content detection, deletion or downgrading. Nevertheless, we argue that additional measures must be considered to actively prevent the issues that stem from attention capturing rather than only mitigating their impact after they have occurred. Presumably, such measures would be political as well as technical, meaning that this socio-technical problematic situation calls for socio-technical solutions. And although the measures may not be associated with immediate research opportunities for the computer science community, we believe that the potential impacts are crucial enough for the community to be fully aware of, and actively involved in, this reflection.
@@ 2 Users in the loop… of recommendation systems
Content recommendation algorithms are a key component of a wide range of applications, including social media, search engines and major Web platforms in general. Through many applications they have changed our lives, helping us to be more efficient, assisting us in daily tasks, or improving our education and information. In a number of other applications however, the truth in not so bright. In the case of social media for instance, they are presented to us as if designed to provide us with content that matches our needs and desires, while what they really seek is to maximize the attention we pay to their hosting platform and advertisements thereof.
Through the training process, recommendation algorithms automatically learn to extract from massive behavioral traces the content’s features that most effectively capture our attention and maximize our continuous engagement with the platform. For instance, they can learn that some categories of topics, such as conflictuality, fear or sexuality, irresistibly attract our attention [10], and thus lean toward recommending these particular categories. They can also learn to select content tailored for a certain user by taking into account the content’s features (topics, source, emotions conveyed…) and its adequacy with the user’s profile (interests, inclinations, past behavior…). This adequacy likely involves many other features that are not transparent since the platforms rarely inform users about how and for which purpose their personalized feed was composed. This is underlined in a study by DeVito [18] who analyzed Facebook’s patents, press releases, and Securities and Exchange Commission filings, to identify “the set of algorithmic values that drive the News Feed”. Some of the features he identified are objective, i.e. they can be observed or measured: friend relationships, explicitly expressed user interests, prior engagement, post age and page relationships. By contrast, other features are up to interpretation and thus raise multiple questions: implicitly expressed user preferences (what are the signals of such implicit expression?), platform priorities (what are they and who decides them?), content quality (what are the quality criterion?).
Finally, it may seem that recommendation algorithms learn to leverage psychological traits and cognitive biases. Yet, it is important to stress that the algorithm does not discover such things as a psychological trait or a cognitive bias itself. Rather, it discovers the features that enable it to exploit what psychologists would describe as a trait or bias. Such criteria are not explicitly formulated, they may not even be explicable nor verifiable. They remain implicit in the models unless a study be carried out a posteriori, that would surface the biases that emerge from the recommendations. This is yet another example where AI techniques without explanations nor feedback are problematic.
@@
A classical self-reinforcing recommendation loop is illustrated in figure 1: 1 The algorithms recommend content to the user. 2 The behavior of the user is captured, possibly partially due to the focus of the platform and the limited choices that the interface offers, and possibly biased due to the fact that these choices may be oriented, again by the interest of the platform and the chosen interface. 3 The algorithms integrate these reactions in future recommendations. As a result, the reactions of the user will reinforce the recommendation and propagation of the attached content.
@@
Of course there are externalities to that loop, that can increase its impact. Smartphones, for example, provide additional means to profile users by tracking their every moves, making recommendation even more efficient and targeted to the point that it competes and sometimes takes over more traditional ways of advertising [65]. Another (detrimental) externality of this loop is that it opens the door to spoofing techniques and other malevolent actors intentionally biasing usage traces to “hijack” recommendation systems. Indeed, as soon as a process is known and documented, it runs the risk of being diverted from its original purpose and manipulated beyond its original objectives. For instance, fake reviews and reactions alter recommendations; black hat techniques of SEO (Search Engine Optimization) such as hidden texts, link farms, cloaking [^6] Cloaking denotes a technique in which the content presented to a search engine crawler is different from that presented to an actual user. It aims at deceiving search engines so they display the page that they would otherwise downgrade or dismiss. Adapted from https://en.wikipedia.org/wiki/Cloaking. or text spinning are disapproved by search engines as they impact the recommendations they make by unduly increasing the ranking of targeted pages or avoiding their downgrading.
As a result, the fact that a few recommendation systems influence a significant fraction of the human population may have a number of detrimental side effects on their users and our societies at large. A first side effect is that recommendation algorithms tend to lock users in an informational space in accordance with their tastes and beliefs, a “filter bubble” [45] that confines them to a “cognitive comfort zone” and activates their confirmation bias as they are faced with information which seems to go towards the same directions or conclusions [55, 34] Eventually, users are no longer confronted with contradiction, debate nor disturbing facts or ideas, and this algorithmic amplification tends to be a powerful driver of the radicalization and polarization of opinions, leading to extremist ideas in some cases [74].
@@ 3 Algorithmic Emotional Governance
Considering the platform’s recommendation loop introduced in section 2, we now want to stress that, directly or indirectly, emotions are a key feature of the selected recommendations. In fact, the whole attention market could be seen as driven by a complex equation involving, at least, emotions, cognitive biases and content recommendation algorithms. This could lead to what we will call here an algorithmic emotional governance merging two concepts: emotional governance [48] which is the informed management of the emotional dynamics of the governed population, and algorithmic governance [54] which is a governance of our societies based on the algorithmic processing of massive data.
@@
Emotions are a powerful attractor of our attention, especially emotions with a high negative valence [61]. As a result, information that arouses anger, fear, indignation, resentment, frustration or disgust is among those that most effectively catch our attention [50, 35]. An explanation is that witnessing others’ negative emotions activates our comparison bias and subjects us to some sort of injunction to take sides, to show our emotional response, and hence publicly demonstrate our “irreproachable morality” [17, 10]. Note that catching attention and increasing user engagement are different things, and although high-arousal negative emotions catch attention more efficiently than other emotions, it remains unclear whether they induce a higher user engagement on social media. In some cases a higher sharing rate of information conveying positive emotions was observed [33, 36]. Nevertheless, in several contributions, researchers showed the overwhelming impact of emotions in argumentation and debates and the means to detect them  [5, 68, 4], and it has also been shown that anger spreads faster on social media than any other emotion [21]. Note that this attraction for negative content can be observed in completely different domains, e.g. in literature where the anti-utopian and dystopian fiction genres became more prominent within the utopian genre [39].
@@
Eventually, nuance, doubt or agnosticism are mechanically made invisible because the low emotional response that they induce simply downgrades their ranking. It is imperative to have an opinion, preferably definite and cleaving. Amplified by digital disinhibition888the feeling of impunity induced by the feeling of anonymity [63], this emotional escalation can lead to outpourings of violence and hatred whose outcome is sometimes tragic as attested by the suicide attempts of teenagers being cyberbullied [56]. Moreover, the full consequences of triggering or regulating emotions on our cognitive functions in general and on memory in particular remain to be studied extensively [49].
@@
We just described the combined effect of emotions, cognitive biases and recommendation algorithms, which is at work whatever the type of content a platform serves. But things get even worse when it comes specifically to false information. False information are frequently meant to arouse strong negative emotions [75], and the combination with cognitive biases and recommendation algorithms provides them with a particularly fertile ground and a formidable cognitive efficiency [2, 38]. Some studies reported that negatively biased fake news enhance users’ willingness to share them [14], and reveal a positive correlation between the virality of fake news and the anger they carry [12]. Another study contended that falsehood spreads “significantly farther, faster, deeper, and more broadly than the truth” on social media [69], which underlines that recommendation of content arousing negative emotions does not only induce local individual reaction: it creates a chain reaction leveraging the network effect of social media to spread that “content–emotion” couple through the acquaintance links. Other studies reported that recommendation algorithms mechanically tend to favor false information conveying divisive ideas, shocking events and negative emotions [22, 23]. This type of content entails a felt injunction to take sides and compulsively spread shocking information rather than appealing to critical thinking, questioning its veracity and verifying its source. And since this information is often relayed by acquaintances, the social proof bias [13] entices users to deem it credible and trustworthy.
@@
Concerningly, the contents we are exposed to leave a trace in our implicit memory: although we cannot recall seeing it, it may impact our choices for several months [15]. Even more concerning is the fact that, due to the negativity bias, negative information leaves a longer memory trace than positive information. Therefore, even when a false information is denied or rectified, there remains a negative feeling that stems from the strong emotional response it triggered in the first place. Repeated again and again, associated with representations of the world that summon conspiracy theories, reinforced under the pressure of filter and emotion bubbles, propelled by the network effect, such information gradually and insidiously undermines our trust in the elites (scholars, experts, journalists, politicians, etc.), entails risks for public health [72, 51], and spurs the emergence of extreme ideas and populism that eventually undermine democracies [74, 1, 30, 23], among other pitfalls.
@@
Finally, let us stress that if “previous studies have shown how personality, values, emotions and vulnerability of users affect their likelihood to propagate misinformation” [22], in this section we only considered an average user without any particular health condition. But we should envisage more complex situations when it comes to users with disabilities or mental disorders e.g. depression, anxiety, compulsive shopping disorder, paranoia, FOMO, FOBO [^9] Fear of Better Options: the inability to choose when faced with a multitude of options. … Let us just mention one specific condition: the attention deficit (AD) disorder. There is evidence that AD symptoms could be worsened by the use of digital media and their attention-grabbing applications, and more importantly that these applications could provoke AD among people without previous record of such a disorder [47]. To the very least, more research is needed in this respect.

### v2

@@ 1 An Unsustainable Attention Market
Since the advent of mass market in the 50’s, media and advertisement providers have relentlessly tried to figure out effective methods to capture our attention and turn it into revenue. During the last two decades, supported by advances in artificial intelligence (AI), major online social media and Web platforms have brought this process of capturing attention to an unprecedented scale. Based almost exclusively on advertising revenues, their business model consists in providing free services that, in return, collect behavioral traces. This data is then used to maximize the impact of advertisements on users by 1 ensuring their mental availability at the time of being shown the advertisement, and 2 ensuring that the message meets their interests, beliefs and moods (i.e. targeted advertising). Based on research in psychology, sociology and neuroscience, several actors including online social media, games and Web platforms have engineered techniques capable of very effectively plundering our “available brain time” (Lewis 2017; Harris 2019). We can distinguish two broad categories of such techniques.
Firstly, some techniques are explicitly designed to leverage cognitive biases as a means to capture attention. For instance, the likes collected after posting content activate the brain’s dopaminergic pathways (involved in the reward system) and tap into our need for social approval, giving “bright dings of pseudo-pleasure” (Lewis 2017); notifications of smartphone applications feed our appetite for novelty and surprise such that it is difficult to resist the urge to check them; the pull-to-refresh mechanism (Lewis 2017), alike slot machines, exploits the variable reward pattern whereby each time we pull down the screen we may get an update or nothing at all; infinite scrolling (of news, posts or videos…) traps us because of our fear of missing out important information (FOMO) to the point that we can hardly break the flow; automatic video chaining replaces a deliberate action to continue watching with a required action to stop watching, and entails a frustrating feeling of incompleteness when stopped; etc. Similarly, some techniques harness dark patterns [^1] The legal definition in California is “A user interface is a dark pattern if the interface has the effect of substantially subverting or impairing user autonomy, decision-making, or choice. A business’s intent in designing the interface is not determinative in whether the user interface is a dark pattern, but a factor to be considered.” CPRA § 7004 (c) (Gray et al. 2023) to manipulate users into taking actions or decisions they wouldn’t take otherwise. This is typically the case when one accepts all notifications of an application without really noticing it, while deactivating notifications would require an additional, less intuitive, series of actions.
@@
Secondly, recent advances in machine learning allow the training of content recommendation algorithms on massive online behavioral data, which Zuboff refers to as “behavioral surplus” (Zuboff 2019). These algorithms learn to recommend content that not only captures attention but also increases user engagement [^2] There are multiple definitions of user engagement. In the context of social media, this typically refers to the fact that a user would interact with a content: e.g. like, comment or repost it. Engagement is usually public in that it leaves public traces on the platform, unlike sheer content consumption that remains private (Robertson et al. 2023).. They discover the content’s key features that help predict whether such content will effectively attract users’ attention, and typically end up selecting content related to conflictuality, fear or sexuality (Bronner 2021). They also learn to exploit humans’ negativity bias (Soroka, Fournier, and Nir 2019; Siegrist and Cvetkovich 2001) and, as a consequence, content conveying high-arousal negative emotions (such as anger, resentment, indignation and disgust) are more likely to be read and eventually shared online than those conveying other emotions (Robertson et al. 2023; Kohout, Kruikemeier, and Bakker 2023). Concerningly enough, false information (a broad term including misinformation and other forms of disinformation) typically relies on such negative emotions as a trick to foster sharing. Finally, recommendation systems may do all this without it being explicit, neither in the features they select nor in the succinct feedback that some of them happen to provide users with [^3] For instance, a recommendation system may tell us “you liked this movie, you may also like this one”. But we don’t know what features were selected to recommend this one: Do they have an actor in common? Did my contacts like both of them? etc..
@@
Since the amount of attention available is both limited and precious, it would be unsustainable for a civilization to waste it massively and with impunity for questionable or futile purposes (Bronner 2021). Today, we might precisely be at that moment: while mental time has become a new oil, we have created an attention economy and subsequent attention markets (Hefti and Heinke 2015; Hendricks and Vestergaard 2019) that, although sustainable from an economic point of view, may be unsustainable from a civilization point of view. From these first references, let us define what the term “attention market” refers to in this article.
@@
The attention market treats attention as a tradable commodity and involves multiple actors: from producers (the end users whose attention is the resource), to content creators whose work is used to capture the attention, brokers who trade and monetize the attention, and consumers (the large Web platforms) who use it for their own purposes such as exposing users to advertisements.
In this article, we propose a discussion aimed at spurring introspection and debate, primarily within the computer science community which designs these technologies, but also more generally within all the communities concerned about societal issues of AI technologies. Our goal is not to provide an additional piece of evidence, for there are many as we will see, but to insist on the interdisciplinary recognition of the situation, of its urgency and of the possible means of action. In line with the Web Science Manifesto (Berendt et al. 2021) calling for interdisciplinary approaches to prepare the future of the Web, we bring together and synthesize the conclusions of more than 70 papers and books from a wide range of disciplines to analyze the practices and drifts of these systems designed to capture attention on a worldwide scale. We make the point that, with the initial commonplace goal of making targeted advertising more effective, the generalization of attention-capturing techniques and their use of negative emotions tends to foster radicalization and polarization, amplify the dissemination of false information, spur the emergence of populism, and eventually put a threat on democracies and human societies in general.
If we look specifically at the computer science community, it appears that, so far, the public and private research has invested large efforts in dealing with some aspects of the problem like radicalization, violent speech and false information. Most of these works rely on post hoc measures (measures taken after the problems have appeared) such as content detection, deletion or downgrading (decreasing the content ranking so that it is less likely to be recommended). Nevertheless, we argue that additional measures must be considered to actively prevent the issues that stem from attention capturing rather than only mitigating their impact once they have occurred. Presumably, such measures would be political and economical as well as technical, meaning that this socio-technical problematic situation calls for socio-technical solutions. Hence this call for an inter-disciplinary approach meant to regulate the attention market.
@@ 2 Users in the Loop… of Recommendation Systems
Content recommendation algorithms are a key component of a wide range of applications, including social media, search engines and major Web platforms in general. Through many applications they have changed our lives, helping us to be more efficient, assisting us in daily tasks, or improving our education and information. In a number of other applications however, the reality in not so bright. In the case of social media for instance, they are presented to us as if designed to provide us with content that matches our needs and desires, while what they really seek is to maximize the attention we pay to their hosting platform and advertisements thereof.
Through the training process, recommendation algorithms automatically learn to extract from massive behavioral traces the content’s features that most effectively capture our attention and maximize our continuous engagement with the platform. Typically, they can learn that some categories of topics, such as conflictuality, fear or sexuality, irresistibly attract our attention (Bronner 2021), and thus lean toward recommending these particular categories. They can also learn to select content tailored for a certain user, by taking into account both the content’s features (topics, source, emotions conveyed…) and its adequacy with the user’s profile (interests, inclinations, past behavior…). This adequacy likely involves many other features that are not transparent since the platforms rarely inform users about how and for which purpose their personalized feed was composed. This is underlined in a study by DeVito (DeVito 2017) who analyzed Facebook’s patents, press releases, and Securities and Exchange Commission filings, to identify “the set of algorithmic values that drive the News Feed”. Some of the features he identified are objective, i.e. they can be observed or measured: friend relationships, explicitly expressed user interests, prior engagement, post age and page relationships. By contrast, other features are up to interpretation and thus raise multiple questions: implicitly expressed user preferences (what are the signals of such implicit expression?), platform priorities (what are they and who decides them?), content quality (what are the quality criterion?).
Finally, it may seem that recommendation algorithms learn to leverage psychological traits and cognitive biases. Yet, it is important to stress that the algorithm does not discover such things as a psychological trait or a cognitive bias itself. Rather, it discovers the features that enable it to exploit what psychologists would describe as a trait or bias. Such criteria are not explicitly formulated, they may not even be explicable nor verifiable. They remain implicit in the models unless a study be carried out a posteriori, that would surface the biases that emerge from the recommendations. This is yet another example where AI techniques, without explanations nor feedback, raise ethical concerns.
@@
A classical self-reinforcing recommendation loop is illustrated in figure 1: 1 The algorithms recommend content to the user. 2 The behavior of the user is captured, possibly partially due to the focus of the platform and the limited choices that the interface offers, and possibly biased due to the fact that these choices may be oriented, again by the interest of the platform and the chosen interface. 3 The algorithms integrate these reactions in future recommendations. As a result, the reactions of the user will reinforce the recommendation and propagation of the content.
@@
Of course there are externalities to that loop, that can increase its impact. Smartphones, for example, provide additional means to profile users by tracking their every moves, making recommendation even more efficient and targeted to the point that it competes and sometimes takes over more traditional ways of advertising (Times 2017). Another (detrimental) externality of this loop is that it opens the door to spoofing techniques and other malevolent actors intentionally biasing usage traces to “hijack” recommendation systems. Indeed, as soon as a process is known and documented, it runs the risk of being diverted from its original purpose and manipulated beyond its original objectives. For instance, fake reviews and reactions alter recommendations; black hat techniques of SEO (Search Engine Optimization) such as hidden texts, link farms, cloaking [^4] Cloaking denotes a technique in which the content presented to a search engine crawler is different from that presented to an actual user. It aims at deceiving search engines so they display the page that they would otherwise downgrade or dismiss. Adapted from https://en.wikipedia.org/wiki/Cloaking. or text spinning are disapproved by search engines as they impact the recommendations they make by unduly increasing the ranking of targeted pages or avoiding their downgrading.
As a result, the fact that a few recommendation systems influence a significant fraction of the human population may have a number of detrimental side effects on their users and our societies at large. A first side effect is that recommendation algorithms tend to lock users in an informational space in accordance with their tastes and beliefs, a “filter bubble” (Pariser 2011) that confines them to a “cognitive comfort zone” and activates their confirmation bias as they are faced with information which seems to go towards the same directions or conclusions (Sasahara et al. 2021; Kitchens, Johnson, and Gray 2020). Eventually, users are no longer confronted with contradiction, debate nor disturbing facts or ideas, and this algorithmic amplification tends to be a powerful driver of the radicalization and polarization of opinions, leading to extremist ideas in some cases (Whittaker et al. 2021).
@@ 3 Algorithmic Emotional Governance
Considering the platform’s recommendation loop introduced in section 2, we now want to stress that, directly or indirectly, emotions are a key feature of the selected recommendations. In fact, the whole attention market could be seen as driven by a complex equation involving, at least, emotions, cognitive biases and content recommendation algorithms. This could lead to what Patino calls an “emocraty, a regime that makes our emotions become performative and invade public space” (Patino 2024). Here we will rather speak of an algorithmic emotional governance merging two concepts: emotional governance (Richards 2007) which is the informed management of the emotional dynamics of the governed population, and algorithmic governance (Rouvroy and Berns 2013) which is a governance of our societies based on the algorithmic processing of massive data.
@@
Emotions are a powerful attractor of our attention, especially emotions with a high negative valence (Soroka, Fournier, and Nir 2019). As a result, information that arouses anger, fear, indignation, resentment, frustration or disgust is among those that most effectively catch our attention (Robertson et al. 2023; Kohout, Kruikemeier, and Bakker 2023). An explanation is that witnessing others’ negative emotions activates our comparison bias and subjects us to some sort of injunction to take sides, to show our emotional response, and hence publicly demonstrate our “irreproachable morality” (Crockett 2017; Bronner 2021). Note that catching attention and increasing user engagement are different things, and although high-arousal negative emotions catch attention more efficiently than other emotions, it remains unclear whether they induce a higher user engagement on social media. In some cases a higher sharing rate of information conveying positive emotions was observed (Keib et al. 2018; Kramer 2012). Nevertheless, in several contributions, researchers showed the overwhelming impact of emotions in argumentation and debates and the means to detect them  (Benlamine et al. 2015; Villata et al. 2018; Basile et al. 2016), and it has also been shown that anger spreads faster on social media than any other emotion (Fan et al. 2014). Note that this attraction for negative content can be observed in completely different domains, e.g. in literature where the anti-utopian and dystopian fiction genres became more prominent within the utopian genre (Mascarell 2020).
@@
Figure 2: Five interfaces of well-known social media to react to a post.
@@
Eventually, nuance, doubt or agnosticism are mechanically made invisible because the low emotional response that they induce simply downgrades their ranking. It is imperative to have an opinion, preferably definite and cleaving. Amplified by digital disinhibition666the feeling of impunity induced by the feeling of anonymity (Suler 2004), this emotional escalation can lead to outpourings of violence and hatred whose outcome is sometimes tragic as attested by the suicide attempts of teenagers being cyberbullied (Schonfeld et al. 2023). Moreover, the full consequences of triggering or regulating emotions on our cognitive functions in general and on memory in particular remain to be studied extensively (Richards and Gross 2000).
@@
We just described the combined effect of emotions, cognitive biases and recommendation algorithms, which is at work whatever the type of content a platform serves. But things get even worse when it comes specifically to false information. False information are frequently meant to arouse strong negative emotions (Zaeem, Li, and Barber 2020), and the combination with cognitive biases and recommendation algorithms provides them with a particularly fertile ground and a formidable cognitive efficiency (Acerbi 2019; Martel, Pennycook, and Rand 2020). Some studies reported that negatively biased fake news enhance users’ willingness to share them (Corbu et al. 2021), and reveal a positive correlation between the virality of fake news and the anger they carry (Chuai and Zhao 2022). Another study contended that falsehood spreads “significantly farther, faster, deeper, and more broadly than the truth” on social media (Vosoughi, Roy, and Aral 2018), which underlines that recommendation of content arousing negative emotions does not only induce local individual reaction: it creates a chain reaction leveraging the network effect of social media to spread that “content–emotion” couple through the acquaintance links. Other studies reported that recommendation algorithms mechanically tend to favor false information conveying divisive ideas, shocking events and negative emotions (Fernández, Bellogín, and Cantador 2021; Fortunato and Pecoraro 2022). This type of content entails a felt injunction to take sides and compulsively spread shocking information rather than appeal to critical thinking, question its veracity and verify its source. And since this information is often relayed by acquaintances, the social proof bias (Cialdini 2008) entices users to deem it credible and trustworthy.
@@
Concerningly, the contents we are exposed to leave a trace in our implicit memory: although we cannot recall seeing it, it may impact our choices for several months (Courbet et al. 2014). Even more concerning is the fact that, due to the negativity bias, negative information leaves a longer memory trace than positive information. Therefore, even when a false information is denied or rectified, there remains a negative feeling that stems from the strong emotional response it triggered in the first place. Repeated again and again, associated with representations of the world that summon conspiracy theories, reinforced under the pressure of filter and emotion bubbles, and propelled by the network effect, such information tends to gradually and insidiously undermine our trust in experts (scholars, journalists, etc.), entails risks for public health (Waszak, Kasprzycka-Waszak, and Kubanek 2018; Rocha et al. 2023), and spurs the emergence of extreme ideas and populism that eventually undermine democracies (Whittaker et al. 2021; Abreu and França 2021; Intelligence 2022; Fortunato and Pecoraro 2022), among other pitfalls.
@@
Finally, let us stress that if “previous studies have shown how personality, values, emotions and vulnerability of users affect their likelihood to propagate misinformation” (Fernández, Bellogín, and Cantador 2021), in this section we only considered an average user without any particular health condition. But we should envisage more complex situations when it comes to users with disabilities or mental disorders e.g. depression, anxiety, compulsive shopping disorder, paranoia, FOMO, FOBO [^7] Fear of Better Options: the inability to choose when faced with a multitude of options. … Let us just mention one specific condition: the attention deficit (AD) disorder. There is evidence that AD symptoms could be worsened by the use of digital media and their attention-grabbing applications, and more importantly that these applications could provoke AD among people without previous record of such a disorder (Ra et al. 2018). To the very least, more research is needed in this respect.